OpenStudy (anonymous):

integral of (3t-2)/(t+1)

jimthompson5910 (jim_thompson5910):

\[\Large \int \left(\frac{3t-2}{t+1}\right) dt\] \[\Large \int \left(\frac{3(u-1)-2}{u}\right) du\] \[\Large \int \left(\frac{3u-3-2}{u}\right) du\] \[\Large \int \left(\frac{3u-5}{u}\right) du\] \[\Large \int \left(\frac{3u}{u}-\frac{5}{u}\right) du\] \[\Large \int \left(3-\frac{5}{u}\right) du\] I'll let you finish up

jimthompson5910 (jim_thompson5910):

I let u = t+1, so t = u-1 Deriving both sides of u = t+1 with respect to t gives du/dt = 1 du = dt

OpenStudy (anonymous):

this helps a lot man i appreciate it. missed class when we learned this...

OpenStudy (anonymous):

then you just substitute back in after you have the final integral?